# GenCeption: Evaluate Multimodal LLMs with Unlabeled Unimodal Data

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) are typically assessed using expensive annotated multimodal benchmarks, which often lag behind the rapidly evolving demands of MLLM evaluation. This paper outlines and validates GenCeption, a novel, annotation-free evaluation method that requires only unimodal data to measure inter-modality semantic coherence and inversely assesses MLLMs’ tendency to hallucinate. This approach eliminates the need for costly data annotation, minimizes the risk of training data contamination, results in slower benchmark saturation, and avoids the illusion of emerging abilities. Inspired by the DrawCeption game, GenCeption begins with a non-textual sample and proceeds through iterative description and generation steps. The semantic drift across iterations is quantified using the GC@$T$ metric. Based on the GenCeption method, we establish the MMECeption benchmark for evaluating Vision LLMs (VLLMs), and compare performance of several popular VLLMs and human annotators. Our empirical results validate GenCeption’s effectiveness, demonstrating strong correlations with established VLLM benchmarks. VLLMs still significantly lack behind human performance and struggle especially with text-intensive tasks.
GenCeption: Evaluate Multimodal LLMs with Unlabeled Unimodal Data
@@ 1 Introduction
Large Language Models (LLMs) demonstrate exceptional abilities in natural language understanding, reasoning, and problem-solving. Multimodal LLMs (MLLMs) enhance these capabilities by incorporating multiple modalities, with the visual modality being predominant Achiam et al. (2023); Liu et al. (2023b); Jiang et al. (2023); Ye et al. (2023). Building on LLMs, MLLMs integrate non-textual modalities, enabling richer interactions and broader applications in real-world scenarios. However, there is a lack of comprehensive evaluation methods to compare different MLLM architectures and training approaches Fu et al. (2023).
In response, the community has developed several MLLM benchmarks, as detailed by Xu et al. (2022); Dai et al. (2023); Wang et al. (2023); Ye et al. (2023); Li et al. (2023b); Zhao et al. (2023). Yet, these benchmarks face common challenges: (1) They predominantly rely on multimodal datasets that demand high-quality annotations, which is costly and restrictive in capturing the evolving capabilities of MLLMs Fu et al. (2023). This has been shown to result in increasing speed in benchmark saturation while contemporary models still struggle on trivial real-world tasks Kiela et al. (2021). (2) MLLM evaluation benchmarks rely on discrete metrics like accuracy that may falsely suggest emergent abilities and do not allow predictable projections of performance improvements from model scaling Schaeffer et al. (2023). (3) The evaluation scores may not reflect true performance on real-world tasks due to potential contamination of MLLM training data by benchmark datasets, as reported for LLM pretraining corpora Dodge et al. (2021); Yang et al. (2023). (4) The content of one modality is often not needed to answer benchmark questions, as the answer can often be inferred from the question or the MLLM’s pretraining knowledge. As a consequence of both (3) and (4), some MLLMs can excel on vision QA benchmarks without even being provided the image that is associated with the question.
@@
Figure 1: An illustration of the $t$-th iteration in the GenCeption evaluation procedure for MLLMs. Using the image modality as an example, the process begins with an existing image $\mathbf{X}^{(0)}$ sourced from a unimodal image dataset for the first iteration ($t$=1). The MLLM provides a detailed description of the image, which is then used by an image generator to produce $\mathbf{X}^{(t)}$.
@@
To address these highlighted challenges, we propose GenCeption, a novel and simple approach for evaluating MLLMs. GenCeption addresses challenge (1) by relying on easily accessible unimodal datasets, which allows for cost-effective and scalable benchmark creation. Relying on unimodal datasets additionally addresses challenge (3) and (4), as it allows to easily use previously unseen datasets for MLLM evaluation, and enforces the relevance of the provided modality for excelling at this task. To tackle challenge (2), GenCeption uses the continuous GC@$T$ metric, providing a more nuanced evaluation compared to discrete metrics, allowing for better projections of performance improvements and avoiding false suggestions of emergent abilities.
@@
GenCeption assesses MLLMs’ ability to consistently maintain semantic coherence across modalities by iteratively generating and describing non-textual samples and measures the continuous GC@$T$ metric. This approach simultaneously evaluates the MLLM’s tendency to hallucinate, as this inversely correlates with semantic coherence. Further, an MLLM’s ability to provide complete yet concise descriptions of non-textual samples measures a diverse range of specialised abilities. For instance, to perform well at describing an image using a limited number of tokens, it is advantageous to be able to reason over people’s emotions and intentions behind their actions, infer the current and preceding weather, count objects, and recognize artistic styles. This list can be extended to various abilities depending on the non-textual modality and the content of the samples used during the GenCeption process.
Our contributions are the following:
@@
  • Proposing GenCeption, an evaluation method that allows for using unlabelled unimodal datasets for MLLM evaluation.
@@
  • Introducing MMECeption, a Vision LLM (VLLM) evaluation benchmark utilizing the GenCeption method. MMECeption uses the images from the MME benchmark Fu et al. (2023), but without their annotated question and answer pairs.
@@ 2 GenCeption
Our approach, GenCeption, is inspired by a multi-player game DrawCeption [^1] https://wikipedia.org/wiki/drawception (a.k.a., Scrawl or Whispernary). In this game, the first player is given an image and describes it verbally to the next player. This player then attempts to recreate the image based on the description. The cycle repeats, often resulting in amusing deviations from the original image. The challenge and objective are to maintain the initial information through iterative transitions between verbal descriptions and drawings. Similarly, a proficient Multimodal Language Model (MLLM), which models multiple modalities such as text and images, should excel at this task. Recognizing that MLLMs can encompass modalities beyond just visual cues, such as audio and graphs, we name our approach GenCeption, covering a broader scope than the visually-centric DrawCeption.
@@ 2.1 Procedure
Unlike existing MLLM benchmarks that rely on multimodal samples, GenCeption is designed to operate on unimodal datasets, significantly streamlining dataset acquisition efforts. For illustrative purposes, we employ the image modality as a representative non-textual modality throughout this exposition. Let’s consider an image dataset $\bm{\mathcal{D}}$ comprising images $\mathbf{X}_{1},\mathbf{X}_{2},\ldots,\mathbf{X}_{N}$, similar to well-established datasets like ImageNet Deng et al. (2009), CIFAR Krizhevsky et al. (2009), and STL Coates et al. (2011). Without loss of generality, any image from $\bm{\mathcal{D}}$ is denoted as $\mathbf{X}$.
@@
GenCeption operates iteratively, from $t=1$ to a pre-defined maximum iteration $t=T$. Each iteration, as depicted in Figure 1, begins with an image $\mathbf{X}^{(t-1)}$ and yields a new image $\mathbf{X}^{(t)}$. The first iteration ($t=1$) starts with the original image $\mathbf{X}^{(0)}$ from $\bm{\mathcal{D}}$. During any given iteration $t$, the MLLM receives a textual prompt $\mathbf{P}_{\text{Desc}}$ (Table 1), instructing the MLLM to generate a comprehensive description $\mathbf{Q}_{t}$ for the input image $\mathbf{X}^{(t-1)}$:
$\mathbf{Q}_{t}:=\text{MLLM}(\mathbf{P}_{\text{Desc}},\mathbf{X}^{(t-1)}).$ (1)
Following this, an image generation prompt $\mathbf{P}_{\text{Gen}}^{(t)}$ is constructed as “Generate an image that fully and precisely reflects this description: <$\mathbf{Q}_{t}$>”. This prompt guides a pretrained image generation model, such as DALL·E Ramesh et al. (2021) or Imagen DeepMind (2023), to create a new image, $\mathbf{X}^{(t)}$:
@@
The textual prompt $\mathbf{P}_{\text{Desc}}$ is intentionally kept short and concise to minimize potential variations in model behaviours due to susceptibility to prompt composition Loya et al. (2023).
@@ 2.2 Metric: GC@$T$
Our primary objective is to measure the semantic divergence of each generated image $\mathbf{X}^{(t)}$ (for $t\in\{1,2,\ldots,T\}$) from the original image $\mathbf{X}^{(0)}$. We use a pretrained image encoder, such as ViT Dosovitskiy et al. (2021), to transform all images, resulting in $T+1$ image embeddings denoted as $\mathbf{z}^{(0)},\mathbf{z}^{(1)},\ldots,\mathbf{z}^{(T)}$, where $\mathbf{z}^{(t)}:=\text{Enc}(\mathbf{X}^{(t)})$. We then compute the cosine similarity between $\mathbf{z}^{(0)}$ and each $\mathbf{z}^{(t)}$ (for $t\in\{1,2,\ldots,T\}$), yielding $T$ similarity scores: $s^{(1)},s^{(2)},\ldots,s^{(T)}$. Here, $s^{(t)}\in[-1.0,1.0]$ approximates the level of semantic drift in the $t$-th iteration of the GenCeption procedure. To quantify the overall speed and magnitude of semantic drift, we calculate the GenCeption score over $T$ iterations, denoted as GC@$T\in[-1.0,1.0]$, as follows:
@@
Input: MLLM to be evaluated, a unimodal dataset $\bm{\mathcal{D}}$:
$\smash{\mathbf{X}_{1}^{(0)},\ldots,\mathbf{X}_{n}^{(0)},\ldots,\mathbf{X}_{N}^{(0)}}$, fixed textual prompt $\mathbf{P}_{\text{Desc}}$,
a sample generator Gen(·), and a sample encoder Enc(·)
@@ 9            Generate description $\mathbf{Q}_{t}$ for $\mathbf{X}_{n}^{(t-1)}$ using (1);
10            Create sample generation prompt $\smash{\mathbf{P}_{\text{Gen}}^{(t)}}$;
11            Generate a new sample $\mathbf{X}_{n}^{(t)}$ according to (2);
12            $s^{(t)}$ := CosineSimilarity($\smash{\mathbf{z}^{(0)}}$, Enc($\smash{\mathbf{X}_{n}^{(t)}}$));
@@
15      Calculate GC@$T$ += $\smash{\sum_{t=1}^{T}(t\cdot s^{(t)})/\sum_{t=1}^{T}t}$; (3)
@@
18return GC@$T$ / N;
Algorithm 1 Calculate GC@$T$ via GenCeption
@@ 3 Experiments
We run several extensive experiments to validate the GenCeption method by comparing the GC@$T$ scores achieved by several VLLMs to the scores they achieve on carefully crafted established benchmarks and to average human performance. Although GenCeption’s design merely requires unimodal image datasets, we employ the same data as used by a recent and well-validated MLLM benchmark, MME Fu et al. (2023). While we discard the annotated question-answer pairs associated with the images in this benchmark, this provides us with the ability (1) to facilitate direct comparison with metrics that include textual QA (question-answering) annotations, and (2) to enable a detailed assessment of MLLM performance across MME’s 14 meticulously crafted sample categories. Attributing this newly created benchmark to the MME dataset and the GenCeption method, we refer to it as the MMECeption benchmark.
We select seven VLLMs – Gemini1.5-Pro Reid et al. (2024), GPT-4o OpenAI (2024) , GPT-4V Achiam et al. (2023), Claude3-Opus Anthropic (2023), LLaVA-7B/13B Liu et al. (2023b) and mPLUG-Owl2 Ye et al. (2023) – based on their superior performance on the OpenCompass multimodal leaderboard OpenCompass (2023), which incorporates a comprehensive set of benchmarks like MME Fu et al. (2023), HallusionBench Liu et al. (2023a), MMStar Chen et al. (2024), SeedBench Li et al. (2023a), and AI2D Kembhavi et al. (2016). We use DALL·E as the default image generation model. To exclude a bias towards MLLMs developed by OpenAI which may have had access to DALL·E generated images during training, we perform an additional evalutaion of all VLLMs on the GC@$1$ score using Imagen2 as an image generation model. We set the temperature parameter to 0 in both the MLLMs and image generators to minimize the stochasticity in model outputs.
As humans are well versed at integrating vision and language modalities, we aim to quantify average human performance on the MMECeption benchmark. As the GenCeption procedure is a labor-intensive and time-consuming task for humans, we randomly select 5 images from each MME category, and by providing human annotators with the same prompts as defined in Table 1, collect results and calculate the GC@$1$ metric. 5 human annotators (3 master students, 1 lecturer, and 1 artist) were recruited to describe one image of each category such that each image in a category is described by a different person to mitigate personal performance differences. The annotators were given 14 weeks to perform this task and were awarded a generous reimbursement of €40 each to ensure sufficient dedication. All annotators were either native English speakers or fluent at a professional level.

### v4

# GenCeption: Evaluate Vision LLMs with Unlabeled Unimodal Data

@@ Abstract
Multimodal Large Language Models (MLLMs) are typically assessed using expensive annotated multimodal benchmarks, which often lag behind the rapidly evolving demands of MLLM evaluation. This paper outlines and validates GenCeption, a novel, annotation-free evaluation method that requires only unimodal data to measure inter-modality semantic coherence and inversely assesses MLLMs’ tendency to hallucinate. This approach eliminates the need for costly data annotation, minimizes the risk of training data contamination, is expected to result in slower benchmark saturation, and avoids the illusion of emerging abilities. Inspired by the DrawCeption game, GenCeption begins with a non-textual sample and proceeds through iterative description and generation steps. The semantic drift across iterations is quantified using the GC@$T$ metric. While GenCeption is principally applicable to MLLMs across various modalities, this paper focuses on its implementation and validation for Vision LLMs (VLLMs). Based on the GenCeption method, we establish the MMECeption benchmark for evaluating VLLMs, and compare the performance of several popular VLLMs and human annotators. Our empirical results validate GenCeption’s effectiveness, demonstrating strong correlations with established VLLM benchmarks. VLLMs still significantly lag behind human performance and struggle especially with text-intensive tasks.
keywords:
multimodal large language model , evaluation , benchmark
journal: Computer Speech & Language
[…]
[L]Preprint published by Computer Speech & Language [https://doi.org/10.1016/j.csl.2025.101785]
[…]
@@ 1 Introduction
Large Language Models (LLMs) demonstrate exceptional abilities in natural language understanding, reasoning, and problem-solving. Multimodal LLMs (MLLMs) enhance these capabilities by incorporating multiple modalities, with the visual modality being predominant and highly commercialized (Achiam et al., 2023; Liu et al., 2023b; Jiang et al., 2023; Ye et al., 2023). Building on LLMs, MLLMs integrate non-textual modalities, enabling richer interactions and broader applications in real-world scenarios. However, there is a lack of comprehensive evaluation methods to compare different MLLM architectures and training approaches (Fu et al., 2023).
In response, the community has developed several MLLM benchmarks, as detailed by Xu et al. (2022); Dai et al. (2023); Wang et al. (2023); Ye et al. (2023); Li et al. (2023c); Zhao et al. (2023). They primarily focus on the visual (i.e., image) and textual input modality due to that VLLMs (Vision LLMs) [^10] Vision Large Language Models (VLLMs) are a specialized subclass of Multimodal Large Language Models (MLLMs) designed to integrate visual and textual modalities for tasks such as image captioning, visual question answering, and multimodal reasoning. While VLLMs are generally capable of processing various visual data types, their most common input is images, owing to the abundance of annotated image-text datasets and the maturity of image processing technologies. are the most widely used and readily available MLLMs on the market. However, these benchmarks face common challenges:
  1. (1)
    They predominantly rely on multimodal datasets that demand high-quality annotations, which is costly and restrictive in capturing the evolving capabilities of MLLMs (Fu et al., 2023). This has been shown to result in increasing speed in benchmark saturation while contemporary models still struggle on trivial real-world tasks (Kiela et al., 2021). Emerging methods like CrossCheckGPT (Sun et al., 2024), designed specifically for MLLM evaluation via cross-system consistency, provide a more relevant, annotation-free alternative. On a broader scope, methods like PRD (Li et al., 2023b) focus on LLM evaluation through peer-based rankings and may be further adapted for MLLM evaluation tasks.
  2. (2)
    MLLM evaluation benchmarks that rely on discrete metrics like accuracy may falsely suggest emergent abilities and do not allow predictable projections of performance improvements from model scaling (Schaeffer et al., 2023).
  3. (3)
    The evaluation scores may not reflect true performance on real-world tasks due to potential contamination of MLLM training data by benchmark datasets, as reported for LLM pretraining corpora (Dodge et al., 2021; Yang et al., 2023).
  4. (4)
    The content of one modality is often not needed to answer benchmark questions, as the answer can often be inferred from the question or the MLLM’s pretraining knowledge.
As a consequence of both (3) and (4), some MLLMs can excel on vision QA benchmarks without even being provided the image that is associated with the question. Existing solutions either only tackle a subset of these challenges, or focus on specific tasks such as image captioning (Lee et al., 2024).
@@
Figure 1: An illustration of the $t$-th iteration in the GenCeption evaluation procedure for VLLMs. Using the image modality as an example, the process begins with an existing image $\mathbf{X}^{(0)}$ sourced from a unimodal image dataset for the first iteration ($t$=1). The VLLM provides a detailed description of the image, which is then used by an image generator to produce $\mathbf{X}^{(t)}$.
@@
We propose GenCeption to address all highlighted challenges involved in the evaluation of task-agnostic MLLMs. GenCeption is designed to be a general evaluation framework that can be applied across modalities. To validate its effectiveness, this paper focuses on Vision LLMs (VLLMs), leveraging the visual modality as an illustrative example. GenCeption addresses challenge (1) by relying on easily accessible unimodal datasets, which allows for cost-effective and scalable benchmark creation. Relying on unimodal datasets additionally addresses challenge (3) and (4), as it allows to easily use previously unseen datasets for MLLM evaluation, and enforces the relevance of the provided modality for excelling at this task. To tackle challenge (2), GenCeption uses the continuous GC@$T$ metric, providing a more nuanced evaluation compared to discrete metrics, allowing for better projections of performance improvements and avoiding the mirage of emergent abilities.
@@
On a high and general level, GenCeption assesses MLLMs’ ability to consistently maintain semantic coherence across modalities by iteratively generating and describing non-textual samples and measures the continuous GC@$T$ metric. This approach simultaneously evaluates the MLLM’s tendency to hallucinate, as this inversely correlates with semantic coherence. Further, an MLLM’s ability to provide complete yet concise descriptions of non-textual samples measures a diverse range of specialized abilities. For instance, to perform well at describing an image using a limited number of tokens, it is advantageous to be able to reason over people’s emotions and intentions behind their actions, infer the current and preceding weather, count objects, and recognize artistic styles. This list can be extended to various abilities depending on the non-textual modality and the content of the samples used during the GenCeption process. The main contributions of this paper are the following:
@@
  • Proposing GenCeption, an evaluation method that principally allows for using unlabeled unimodal datasets for MLLM evaluation.
@@
  • Introducing MMECeption, a Vision LLM (VLLM) evaluation benchmark utilizing the GenCeption method. MMECeption uses the images from the MME benchmark (Fu et al., 2023), but without their annotated question and answer pairs.
@@ 2 GenCeption
Our approach, GenCeption, is inspired by a multi-player game DrawCeption [^11] https://wikipedia.org/wiki/drawception (a.k.a., Scrawl or Whispernary). In this game, the first player is given an image and describes it verbally to the next player. This player then attempts to recreate the image based on the description. The cycle repeats, often resulting in amusing deviations from the original image. The challenge and objective are to maintain the initial information through iterative transitions between verbal descriptions and drawings. Similarly, a proficient Multimodal Language Model (MLLM), which models multiple modalities such as text and images, should excel at this task. Recognizing that MLLMs can encompass modalities beyond just visual cues, such as audio and graphs, we name our approach GenCeption, covering a broader scope than the visually-centric DrawCeption. For the sake of clarity and alignment with our experiments, we will focus on VLLMs in the remainder of this section to walk through the GenCeption approach.
@@ 2.1 Procedure
Unlike existing MLLM benchmarks (often focused on VLLMs) that rely on multimodal samples, GenCeption is designed to operate on unimodal datasets, significantly streamlining dataset acquisition efforts. For illustrative purposes, we employ the image modality as a representative non-textual modality throughout this exposition. Let us consider an image dataset $\bm{\mathcal{D}}$ comprising images $\mathbf{X}_{1},\mathbf{X}_{2},\ldots,\mathbf{X}_{N}$, similar to well-established datasets like ImageNet (Deng et al., 2009), CIFAR (Krizhevsky et al., 2009), and STL (Coates et al., 2011). Without loss of generality, any image from $\bm{\mathcal{D}}$ is denoted as $\mathbf{X}$.
@@
GenCeption operates iteratively, from $t=1$ to a pre-defined maximum iteration $t=T$. Each iteration, as depicted in Figure 1, begins with an image $\mathbf{X}^{(t-1)}$ and yields a new image $\mathbf{X}^{(t)}$. The first iteration ($t=1$) starts with the original image $\mathbf{X}^{(0)}$ from $\bm{\mathcal{D}}$. During any given iteration $t$, the VLLM receives a textual prompt $\mathbf{P}_{\text{Desc}}$ (Table 1), instructing the VLLM to generate a comprehensive description $\mathbf{Q}_{t}$ for the input image $\mathbf{X}^{(t-1)}$:
[…]
Following this, an image generation prompt $\mathbf{P}_{\text{Gen}}^{(t)}$ is constructed as “Generate an image that fully and precisely reflects this description: $<\mathbf{Q}_{t}>$”. This prompt guides a pretrained image generation model, such as DALL·E (Ramesh et al., 2021) or Imagen (DeepMind, 2023), to create a new image, $\mathbf{X}^{(t)}$:
@@
The textual prompt $\mathbf{P}_{\text{Desc}}$ is intentionally kept short and concise to minimize potential variations in model behaviours due to susceptibility to prompt composition (Loya et al., 2023).
@@ 2.2 Metric: GC@$T$
Our primary objective is to measure the semantic divergence of each generated image $\mathbf{X}^{(t)}$ (for $t\in\{1,2,\ldots,T\}$) from the original image $\mathbf{X}^{(0)}$. We use a pretrained image encoder, such as ViT (Dosovitskiy et al., 2021), to transform all images, resulting in $T+1$ image embeddings denoted as $\mathbf{z}^{(0)},\mathbf{z}^{(1)},\ldots,\mathbf{z}^{(T)}$, where $\mathbf{z}^{(t)}:=\text{Enc}(\mathbf{X}^{(t)})$. We then compute the cosine similarity between $\mathbf{z}^{(0)}$ and each $\mathbf{z}^{(t)}$ (for $t\in\{1,2,\ldots,T\}$), yielding $T$ similarity scores: $s^{(1)},s^{(2)},\ldots,s^{(T)}$. Here, $s^{(t)}\in[-1.0,1.0]$ approximates the level of semantic drift in the $t$-th iteration of the GenCeption procedure. To quantify the overall speed and magnitude of semantic drift, we calculate the GenCeption score over $T$ iterations, denoted as GC@$T\in[-1.0,1.0]$, as follows:
@@
Input: VLLM to be evaluated, a unimodal dataset $\bm{\mathcal{D}}$: $\small\smash{\mathbf{X}_{1}^{(0)},\ldots,\mathbf{X}_{n}^{(0)},\ldots,\mathbf{X}_{N}^{(0)}}$, fixed textual prompt $\mathbf{P}_{\text{Desc}}$, a sample generator Gen(·), and a sample encoder Enc(·)
@@ 9            Generate description $\mathbf{Q}_{t}$ for $\mathbf{X}_{n}^{(t-1)}$ using (1);
10            Create $\small\smash{\mathbf{P}_{\text{Gen}}^{(t)}}$ using $\mathbf{Q}_{t}$;
11            Generate $\mathbf{X}_{n}^{(t)}$ according to (2);
12            $s^{(t)}$ := CosineSimilarity($\small\smash{\mathbf{z}^{(0)}}$, Enc($\small\smash{\mathbf{X}_{n}^{(t)}}$));
@@ 11            Generate $\mathbf{X}_{n}^{(t)}$ according to (2);
15      Calculate GC@$T$ += $\small\smash{\sum_{t=1}^{T}(t\cdot s^{(t)})/\sum_{t=1}^{T}t}$; (3)
@@
return GC@$T$ / N;
Algorithm 1 Calculate GC@$T$ via GenCeption for a specific VLLM under evaluation
@@ 3 Experiments
We run several extensive experiments to validate the GenCeption method by comparing the GC@$T$ scores achieved by several VLLMs to the scores they achieve on carefully crafted established benchmarks and to average human performance. Although GenCeption’s design merely requires unimodal image datasets, we employ the same data as used by a recent and well-validated MLLM benchmark, MME (Fu et al., 2023). While we discard the annotated question-answer pairs associated with the images in this benchmark, this provides us with the ability (1) to facilitate direct comparison with metrics that include textual QA (question-answering) annotations, and (2) to enable a detailed assessment of MLLM performance across MME’s 14 meticulously crafted sample categories. Attributing this newly created benchmark to the MME dataset and the GenCeption method, we refer to it as the MMECeption benchmark.
[…]
As humans are well versed at integrating vision and language modalities, we aim to quantify average human performance on the MMECeption benchmark. As the GenCeption procedure is a labor-intensive and time-consuming task for humans, we randomly select 5 images from each MME category, and by providing human annotators with the same prompts as defined in Table 1, collect results and calculate the GC@$1$ metric. Five human annotators (3 master students, 1 lecturer, and 1 artist) were recruited to describe one image of each category such that each image in a category is described by a different person to mitigate personal performance differences. The annotators were given 14 weeks to perform this task and were awarded a generous reimbursement of €40 each to ensure sufficient dedication. All annotators were either native English speakers or fluent at a professional level.
3.1 Quantitative results
We partition the 14 MME categories into two groups based on content type: visual-intensive (10 categories: existence, count, position, color, poster, celebrity, scene, landmark, artwork, and commonsense reasoning) and textual-intensive (4 categories: code reasoning, numerical calculation, text translation, and OCR). GC@$3$ scores on the MMECeption benchmark and accuracy on the original MME benchmark are reported per category and as aggregations in Table 2. Additionally, we include the scores and ranks of all evaluated VLLMs on the OpenCompass (OpenCompass, 2023), MME (Fu et al., 2023), HallusionBench (Liu et al., 2023a), MMStar (Chen et al., 2024), SeedBench (Li et al., 2023a), and AI2D (Kembhavi et al., 2016) leaderboards. Notably, Gemini1.5-Pro leads our rankings, followed by GPT-4o, GPT-4v, Claude3-Opus, mPLUG-Owl2, and LLaVA-13B/7B. The GenCeption method shows robustness to the similarity metric used, as the overall ranking remains identical when using cosine similarity or FID distance for calculating GC@$T$ scores.
GC@1 GC@3 MME HallusionBench
GC@1 (essentially the same as  $s^{(1)}$) - 0.99 0.46 0.92
GC@3 0.99 - 0.52 0.93
$s^{(3)}$ (GC@3 without weighting by $t$) 0.98 0.99 0.46 0.94
CrossCheckGPT (Image-to-text) 0.96 0.94 0.42 0.97
Table 3: Correlation matrix comparing GC@$1$, GC@$3$, $s^{(3)}$ (GC@$3$ without temporal weighting), and CrossCheckGPT with established benchmarks MME and HallusionBench.